OpenStudy (anonymous):

8-5). Perform the following multiplication. The answer should be in simplified form for radical expressions. sqrt(6) sqrt(3)

OpenStudy (anonymous):

sqrt(6) * sqrt(3) = sqrt(18) =sqrt(9*2) = sqrt(9) * sqrt(2) = 3sqrt(2)